### v2

# Gaussian deconvolution and the lace expansion

Abstract
We give conditions on a real-valued function $F$ on $\mathbb{Z}^{d}$, for $d>2$, which ensure that the solution $G$ to the convolution equation $(F*G)(x)=\delta_{0,x}$ has Gaussian decay $|x|^{-(d-2)}$ for large $|x|$. Precursors of our results were obtained in the 2000s, using intricate Fourier analysis. In 2022, a very simple deconvolution theorem was proved, but its applicability was limited. We extend the 2022 theorem to remove its limitations while maintaining its simplicity—our main tools are Hölder’s inequality and basic Fourier theory in $L^{p}$ space. Our motivation comes from critical phenomena in equilibrium statistical mechanics, where the convolution equation is provided by the lace expansion and the deconvolution $G$ is a critical two-point function. Our results significantly simplify existing proofs of critical $|x|^{-(d-2)}$ decay in high dimensions for self-avoiding walk, Ising and $\varphi^{4}$ models, percolation, and lattice trees and lattice animals. We also improve previous error estimates.
1 Gaussian deconvolution
1.1 Introduction
Since its introduction by Brydges and Spencer in 1984, the lace expansion has been extended and applied to prove detailed results concerning mean-field critical behaviour of statistical mechanical models on $\mathbb{Z}^{d}$ above their upper critical dimensions. These results include self-avoiding walk and spin systems (Ising and $\varphi^{4}$) in dimensions $d>4$ (e.g., [6, 5, 19, 30, 31]), percolation for $d>6$ [17, 20, 12], and lattice trees and lattice animals (LTLA) for $d>8$ [18, 13]. An aspect of this work that is central to many applications is a proof that the critical two-point function for each of the mentioned models has Gaussian decay $|x|^{-(d-2)}$ at large distance. A proof of this $x$-space decay was not discovered until suitable deconvolution theorems were proved in the 2000s [15, 16]. This followed two decades of work in which the Fourier transform of the critical two-point function was proved to behave as $|k|^{-2}$ for small frequencies $k$. Knowledge of $|k|^{-2}$ behaviour is sufficient for much progress, including proving the triangle condition [2] for percolation. However, [27, Example 1.6.2] shows that it does not imply $|x|^{-(d-2)}$ decay. The $|x|^{-(d-2)}$ decay has been employed, e.g., in percolation theory to construct and analyse the incipient infinite cluster [21, 23], to study arm exponents [8, 24], and to study finite-size scaling [1, 7, 22] where in particular the near-critical decay $|x|^{-(d-2)}e^{-c(p_{c}-p)^{1/2}|x|}$ of the two-point function has been established and exploited in [22].
In 2018, a simpler proof of an $|x|^{-(d-2)}$ upper bound for the weakly self-avoiding walk in dimensions $d>4$ was presented in [4], using Banach algebras and convolution estimates and without using Fourier transforms. A different and strikingly simple proof of $|x|^{-(d-2)}$ decay for weakly self-avoiding walk in dimensions $d>4$ appeared very recently in [33], where an asymptotic formula was obtained by using the Fourier transform in an elementary way (inspired by Kotani’s Theorem [32]). However, in [33] obstacles were identified to extending the proof to other models including percolation and LTLA. In this paper, we remove the obstacles while maintaining the simplicity of the approach of [33], and we extend it so that its realm of applications includes all models for which $|x|^{-(d-2)}$ decay has been proved via the lace expansion, namely self-avoiding walk, Ising and $\varphi^{4}$ spin models, percolation, LTLA.
This extension provides a new proof of Hara’s Gaussian Lemma [15]. Our method involves an elementary but judicious application of Hölder’s inequality in the context of Fourier analysis for weak derivatives in $L^{p}$ spaces (we recall the necessary theory in Appendix A). We also obtain improved error bounds compared to [15], using a fractional derivative analysis.
More generally, our work provides conditions on a function on $\mathbb{Z}^{d}$ so that its deconvolution has $|x|^{-(d-2)}$ decay at large distance. In applications, the verification of the assumed conditions uses the model-dependent diagrammatic estimates common in lace expansion analysis; we do not discuss that further as the details can be found in [15, 16, 30, 5] and we have nothing new to say in that regard.
In [33], it was also noted that further work would be needed in order to apply the method to spread-out models, which are used to analyse dimensions only slightly above the upper critical dimension. In a companion paper [26], we adapt our deconvolution theorem to spread-out models and use it to significantly simply the proof, compared to [16], of $|x|^{-(d-2)}$ decay of critical two-point functions for spread-out versions of self-avoiding walk, Ising model, percolation, and LTLA above their upper critical dimensions.
Notation: We write $a\vee b=\max\{a,b\}$ and $a\wedge b=\min\{a,b\}$. We write $f=O(g)$ or $f\lesssim g$ to mean there exists a constant $C>0$ such that $\lvert f(x)\rvert\leq C\lvert g(x)\rvert$, and $f=o(g)$ for $\lim f/g=0$. To avoid dividing by zero, with $|x|$ the Euclidean norm of $x\in\mathbb{R}^{d}$, we define
$|\!|\!|x|\!|\!|=\max\{|x|,1\}.$ (1.1)
Note that (1.1) does not define a norm on $\mathbb{R}^{d}$.
1.2 Gaussian deconvolution theorem
The convolution of two absolutely summable functions $f,g:\mathbb{Z}^{d}\to\mathbb{R}$ is the function $f*g:\mathbb{Z}^{d}\to\mathbb{R}$ defined by
$(f*g)(x)=\sum_{y\in\mathbb{Z}^{d}}f(y)g(x-y).$ (1.2)
Let $d>2$, let $D:\mathbb{Z}^{d}\to[0,1)$ be defined by $D(x)=\frac{1}{2d}\mathds{1}_{|x|=1}$, and let $\mu\in(0,1]$. The lattice Green function $C_{\mu}:\mathbb{Z}^{d}\to\mathbb{R}$ is the solution, vanishing at infinity, of
$(\delta-\mu D)*C_{\mu}=\delta,$ (1.3)
where $\delta$ is the Kronecker delta $\delta(x)=\delta_{0,x}$. In other words, $C_{\mu}$ is the deconvolution of the operator $\delta-\mu D$. In particular, $C_{1}$ is the deconvolution of minus the discrete Laplacian, and $C_{1}(x)$ is the expected number of visits to $x$ by a simple random walk started from the origin. The solution to (1.3) can be obtained using the (inverse) Fourier transform
$\hat{f}(k)=\sum_{x\in\mathbb{Z}^{d}}f(x)e^{ik\cdot x}\quad(k\in\mathbb{T}^{d}),\qquad f(x)=\int_{\mathbb{T}^{d}}\hat{f}(k)e^{-ik\cdot x}\frac{dk}{(2\pi)^{d}}\quad(x\in\mathbb{Z}^{d}),$ (1.4)
where $\mathbb{T}^{d}=(\mathbb{R}/2\pi\mathbb{Z})^{d}$ is the continuum torus. We identify $\mathbb{T}^{d}$ with $(-\pi,\pi]^{d}\subset\mathbb{R}^{d}$. In particular, in dimensions $d>2$, $C_{\mu}$ is given by the absolutely convergent integral
$C_{\mu}(x)=\int_{\mathbb{T}^{d}}\frac{e^{-ik\cdot x}}{1-\mu\hat{D}(k)}\frac{dk}{(2\pi)^{d}},\qquad\hat{D}(k)=d^{-1}\sum_{j=1}^{d}\cos k_{j}.$ (1.5)
When $\mu\in(0,1)$, $C_{\mu}$ decays exponentially (detailed asymptotics are given in [29]), whereas for $\mu=1$ the large-$x$ decay of $C_{1}(x)$ is well-known (e.g., [25]) to be given by
$C_{1}(x)=\frac{a_{d}}{|\!|\!|x|\!|\!|^{d-2}}+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d}}\right),\qquad a_{d}=\frac{d\Gamma(\frac{d-2}{2})}{2\pi^{d/2}}.$ (1.6)
We do not give a new proof of (1.6). Although $C_{1}$ is not summable, it is nevertheless a standard fact that (1.5) does give a solution to (1.3) when $\mu=1$.
Motivated by applications to statistical mechanics based on the lace expansion, given a function $F:\mathbb{Z}^{d}\to\mathbb{R}$ we wish to understand the large-$x$ decay of the solution $G$ to the equation
$F*G=\delta,$ (1.7)
given by the Fourier integral
$G(x)=\int_{\mathbb{T}^{d}}\frac{e^{-ik\cdot x}}{\hat{F}(k)}\frac{dk}{(2\pi)^{d}}.$ (1.8)
We are interested in functions $F$ satisfying the following assumptions, which in particular imply the absolute convergence of the above integral.
Assumption 1.1.
We assume that $F$ is a $\mathbb{Z}^{d}$-symmetric function (invariant under reflection in coordinate hyperplanes and rotation by $\pi/2$) for which there are $K_{1},K_{2}>0$ and $\rho>\frac{d-8}{2}\vee 0$ such that, for all $x\in\mathbb{Z}^{d}$ and $k\in\mathbb{T}^{d}$,
$\lvert F(x)\rvert$ $\leq\frac{K_{1}}{|\!|\!|x|\!|\!|^{d+2+\rho}},\qquad\hat{F}(0)\geq 0,\qquad\hat{F}(k)-\hat{F}(0)\geq K_{2}\lvert k\rvert^{2}.$ (1.9)
By Taylor’s Theorem and Assumption 1.1,
$\kappa_{0}=-\sum_{x\in\mathbb{Z}^{d}}|x|^{2}F(x)\in(0,\infty),$ (1.10)
where the equality defines the constant $\kappa_{0}$. Assumption 1.1 is satisfied by $F=\delta-\mu D$ for $\mu\in(0,1]$ for any $\rho>0$. In particular, the infrared bound (lower bound on $\hat{F}(k)-\hat{F}(0)$) in this case is simply
$\mu\hat{D}(0)-\mu\hat{D}(k)=\frac{\mu}{d}\sum_{j=1}^{d}(1-\cos k_{j})\geq\frac{2\mu}{\pi^{2}d}\lvert k\rvert^{2}\qquad(k\in\mathbb{T}^{d}).$ (1.11)
We do not assume that $F(x)\leq 0$ for all $x\neq 0$, unlike analogues of our results such as [34]. This is important for applications to the lace expansion, where positive values of $F$ do occur.
Let
$\lambda$ $=\frac{1}{\hat{F}(0)-\sum_{x\in\mathbb{Z}^{d}}\lvert x\rvert^{2}F(x)}=\frac{1}{\hat{F}(0)+\kappa_{0}},\qquad\mu=1-\lambda\hat{F}(0).$ (1.12)
By hypothesis, $\lambda>0$, so $\mu\leq 1$. Also, $\lambda\hat{F}(0)<1$ by definition of $\lambda$, so in fact $\mu\in(0,1]$. For the critical case of $\hat{F}(0)=0$, we have $\mu=1$ and $\lambda=\kappa_{0}^{-1}$.
The following is our general Gaussian deconvolution theorem. Its proof is given in Section 2.
Theorem 1.2 (Gaussian deconvolution).
Let $d>2$, and let $F$ satisfy Assumption 1.1. Then the solution $G$ to (1.7), given by the Fourier integral (1.8), satisfies
$G(x)=\lambda C_{\mu}(x)+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d-2+s}}\right)$ (1.13)
for any power $s$ which obeys
$s<\rho\wedge 2\qquad$ $(2<d\leq 8),$ (1.14)
$s<(\rho-\tfrac{d-8}{2})\wedge 2\qquad$ $(d>8).$ (1.15)
The error estimate in (1.13) depends only on $d,K_{1},K_{2},\rho,s$.
In the critical case, using $\mu=1$ and the large-$x$ decay of $C_{1}(x)$ in (1.6), the conclusion (1.13) becomes, with $\kappa_{0}$ given by (1.10) and the same restriction on $s$,
$G(x)=\frac{a_{d}}{\kappa_{0}|\!|\!|x|\!|\!|^{d-2}}+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d-2+s}}\right).$ (1.16)
When $\mu\in(0,1)$, which occurs in the subcritical case $\hat{F}(0)>0$, Theorem 1.2 remains useful in the sense that, although $C_{\mu}$ decays exponentially (its detailed asymptotics are given in [29]), we nevertheless obtain an upper bound $G(x)\leq O(|\!|\!|x|\!|\!|^{-(d-2)})$ which is uniform no matter how close to critical $F$ is.
Theorem 1.2 was first proved (in the critical case) by Hara [15] using intricate Fourier analysis, with error exponent [^1] A “slightly lengthy” unpublished improvement of the error term to the presumably optimal $s=\rho\wedge 2$ is mentioned in [15, (1.39)]. $s=(\rho\wedge 2)/d$ and without the assumption that $\rho>\frac{d-8}{2}$. This assumption, which we need for the reason indicated at (2.20), is satisfied for all known applications (see Section 1.3). Our proof of Theorem 1.2 is completely different from the proof in [15] and is much simpler. For the weaker statement that in the critical case (with $\rho>\frac{d-8}{2}\vee 0$ as in Assumption 1.1)
$G(x)=\frac{a_{d}}{\kappa_{0}|\!|\!|x|\!|\!|^{d-2}}+o\left(\frac{1}{|\!|\!|x|\!|\!|^{d-2+s}}\right)\ \text{as $\lvert x\rvert\to\infty$ with}\ s=\begin{cases}0&(\rho\leq 1+(\frac{d-8}{2}\vee 0))\\ 1&(\rho>1+(\frac{d-8}{2}\vee 0)),\end{cases}$ (1.17)
which is sufficient for applications, our proof is remarkably short and simple. The improved error bound in Theorem 1.2 requires more effort, but remains simple conceptually.
Applications to the Ising model, percolation and LTLA involve an inhomogeneous convolution equation of the form
$F*H=g.$ (1.18)
Suppose that $g$ is $\mathbb{Z}^{d}$-symmetric and satisfies
$\lvert g(x)\rvert\leq\frac{K_{3}}{|\!|\!|x|\!|\!|^{d+\rho\wedge 2}}$ (1.19)
for some $K_{3}>0$. We are interested in solutions to (1.18) given by the Fourier integral
$H(x)=\int_{\mathbb{T}^{d}}\frac{\hat{g}(k)e^{-ik\cdot x}}{\hat{F}(k)}\frac{dk}{(2\pi)^{d}}.$ (1.20)
The following simple observation was noted in Hara [15] (with $s=(\rho\wedge 2)/d))$.
Corollary 1.3.
Let $d>2$, and let $F$ satisfy Assumption 1.1 with $\hat{F}(0)=0$ (critical case) and with $g$ obeying (1.19). Then the solution $H$ to (1.18), given by the Fourier integral (1.20), satisfies
$H(x)=\frac{a_{d}\sum_{y}g(y)}{\kappa_{0}|\!|\!|x|\!|\!|^{d-2}}+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d-2+s}}\right)$ (1.21)
with $s$ as in (1.14)–(1.15). The constant in the error term depends only on $d,K_{1},K_{2},K_{3},\rho,s$.
Proof.
Let $G$ be the solution to $F*G=\delta$ given by Theorem 1.2. By (1.20), we have $H=G*g$. The desired result then follows from the asymptotic formula for $G$ in (1.16) and the elementary convolution estimate [16, Proposition 1.7(ii)]. ∎
1.3 Application to statistical mechanical models
For many statistical mechanical models in the high-dimensional setting, the lace expansion provides convolution equations for critical and subcritical two-point functions. Our results can be applied to obtain their $x$-space asymptotics. For nearest-neighbour models, Assumption 1.1 has been verified with
[…]
(For the Ising model, percolation and LTLA, (1.19) is also verified in the above references.) The restriction $\rho>\frac{d-8}{2}\vee 0$ is satisfied in all cases. See also [31] for $\varphi^{4}$. The above references all make use of Hara’s Gaussian Lemma [15, Section 1.2.2]. Our results provide a simpler and more conceptual alternative. In addition, our error bounds in Theorem 1.2 translate directly into the same error bounds for the critical two-point functions of all the above models, which improves previous results. For example, for the critical two-point function of the nearest-neighbour strictly self-avoiding walk in dimensions $d\geq 5$, the use of Theorem 1.2 to replace the Gaussian Lemma in [15] gives
$G_{z_{c}}(x)=\frac{a_{d}}{\kappa_{0}|\!|\!|x|\!|\!|^{d-2}}+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d-\varepsilon}}\right)$ (1.23)
for any $\varepsilon>0$. Similarly, for the critical two-point function of nearest-neighbour percolation in dimensions $d\geq 11$, combined with [12, 15], Corollary 1.3 gives
$H_{z_{c}}(x)=\frac{A_{z_{c}}}{|\!|\!|x|\!|\!|^{d-2}}+O\left(\frac{1}{|\!|\!|x|\!|\!|^{d-\varepsilon}}\right)$ (1.24)
for any $\varepsilon>0$, with an explicit amplitude $A_{z_{c}}$. The errors computed in [15] instead involve $\varepsilon=2-2/d$.
2 Proof of Gaussian deconvolution Theorem 1.2
2.1 Outline of proof
The structure of the proof of Theorem 1.2 is very simple.
Isolation of leading term. Suppose $F$ satisfies Assumption 1.1. For $\mu\in(0,1]$, let $A_{\mu}=\delta-\mu D$, so that $A_{\mu}*C_{\mu}=\delta$ by (1.3). For any $\lambda\in\mathbb{R}$, we write
$G$ $=\lambda C_{\mu}+\delta*G-\lambda C_{\mu}*\delta$
$=\lambda C_{\mu}+(C_{\mu}*A_{\mu})*G-\lambda C_{\mu}*(F*G)$
$=\lambda C_{\mu}+C_{\mu}*E_{\lambda,\mu}*G,$ (2.1)
with $E_{\lambda,\mu}=A_{\mu}-\lambda F$. This isolates the leading term $\lambda C_{\mu}$:
$G$ $=\lambda C_{\mu}+f,\qquad f=f_{\lambda,\mu}=C_{\mu}*E_{\lambda,\mu}*G.$ (2.2)
The first moment of $E$ vanishes by our symmetry assumptions, and our choice of $\mu$ and $\lambda$ has been made to ensure that
$\sum_{x\in\mathbb{Z}^{d}}E_{\lambda,\mu}(x)=\sum_{x\in\mathbb{Z}^{d}}\lvert x\rvert^{2}E_{\lambda,\mu}(x)=0.$ (2.3)
Indeed, (2.3) is a system of two linear equations in $\lambda,\mu$ whose solution is given by (1.12).
Error term with integer power. Consider Theorem 1.2 in the critical case, where the Fourier transforms $\hat{G}(k)$ and $\hat{C}(k)$ both behave like $|k|^{-2}$. As we will show, the vanishing moments of $E$ in (2.3) imply that its Fourier transform is bounded by $|k|^{2+\sigma}$ for any $0<\sigma<\rho\wedge 2$, so that $\hat{f}(k)=\hat{C}(k)\hat{E}(k)\hat{G}(k)$ is of order $|k|^{\sigma-2}$, which is less singular than the main term $\hat{C}(k)$. Roughly speaking, this suggests that a derivative of $\hat{f}$ of order $a$ will have singularity $|k|^{\sigma-2-a}$, which is integrable provided that $a<d-2+\sigma$. For example, if $\sigma\in(1,2)$ then $a=d-1$ is permitted.
As we explain later in more detail, an elementary result from Fourier analysis states that if $a$ is an integer and the derivatives $\nabla^{\alpha}\hat{f}(k)$ are in $L^{1}(\mathbb{T}^{d})$ for $|\alpha|\leq a$, then $f(x)$ decays at least as $o(|x|^{-a})$. This allows for a proof of Theorem 1.2 with integer-power error estimate, by taking $a=d-2$ or $a=d-1$ derivatives of $\hat{f}$.
An important consideration is that the calculation of $\nabla^{\alpha}\hat{f}$ involves calculating $\nabla^{\alpha}\hat{F}$. If we are to take $\lvert\alpha\rvert=d-2$ derivatives, since the sum $\sum_{x\in\mathbb{Z}^{d}}\lvert x\rvert^{d-2}/\lvert x\rvert^{d+2+\rho}$ diverges when $d-2\geq 2+\rho$, the function $x^{\alpha}F(x)$ may not be summable and hence we are not able to conclude that $\nabla^{\alpha}\hat{F}$ exists. This potentially dangerous situation arises for percolation and LTLA, and was pointed out in [33] as an obstacle. We overcome this obstacle using weak derivatives, which allows us to make sense of $\nabla^{\alpha}\hat{F}$ for $\lvert\alpha\rvert<\frac{1}{2}d+2+\rho$ (see (2.20)); it is here that our restriction $\rho>\frac{d-8}{2}$ arises.
Improved error term. In Section 2.3 we will prove that for non-integer $a>0$, if the weak derivative $\nabla^{\alpha}\hat{f}$ is in $L^{1}(\mathbb{T}^{d})$ for all multi-indices $\alpha$ with $\lvert\alpha\rvert=\lfloor a\rfloor$, and if in addition there exists $\eta\in(a-\lfloor a\rfloor,1)$ such that (with $\tilde{u}=(u,0,\ldots,0)\in\mathbb{R}^{d}$)
$\big{\lVert}\nabla^{\alpha}[\hat{f}(\cdot+\tilde{u})-\hat{f}(\cdot-\tilde{u})]\big{\rVert}_{L^{1}(\mathbb{T}^{d})}\lesssim u^{\eta}\qquad(0\leq u\leq 1)$ (2.4)
for all $\lvert\alpha\rvert=\lfloor a\rfloor$, then $f(x)$ decays at least as fast as $|x|^{-a}$. We apply this to obtain the error estimate stated in Theorem 1.2.
2.2 Proof of Theorem 1.2 with integer-power error estimate
We first prove a version of Theorem 1.2 with an integer power of $|\!|\!|x|\!|\!|$ in the error term, which is in fact sufficient for applications. The key ideas in the paper are present already here. In Section 2.3 we improve the error estimate using fractional derivatives and complete the proof of Theorem 1.2.
We assume that $\rho>\frac{d-8}{2}\vee 0$. Let
$n_{d}=d-2+s\quad\text{with}\quad s=\begin{cases}0&(\rho\leq 1+(\frac{d-8}{2}\vee 0))\\ 1&(\rho>1+(\frac{d-8}{2}\vee 0)),\end{cases}$ (2.5)
i.e., we pick the largest integer $s$ that satisfies (1.14)–(1.15). By definition, and since $\rho>\frac{d-8}{2}\vee 0$,
$n_{d}<(d-2+\rho\wedge 2)\wedge(\tfrac{1}{2}d+2+\rho).$ (2.6)
Theorem 2.1.
Let $d>2$, and let $F$ satisfiy Assumption 1.1. Then the solution $G$ to (1.7), given by the Fourier integral (1.8), satisfies
$\qquad~{}G(x)=\lambda C_{\mu}(x)+o\left(\frac{1}{\lvert x\rvert^{n_{d}}}\right)\quad\text{as $\lvert x\rvert\to\infty$}.$ (2.7)
2.2.1 Fourier analysis
Let $F$ satisfy Assumption 1.1. As in (2.2), we write
$G$ $=\lambda C_{\mu}+f,\qquad f=f_{\lambda,\mu}=C_{\mu}*E_{\lambda,\mu}*G,$ (2.8)
where $\lambda$ and $\mu$ have been chosen as in (1.12) in order to ensure that $E_{\lambda,\mu}$ has vanishing zeroth and second moments. The following lemma translates this good moment behaviour into good bounds on $\hat{E}_{\lambda,\mu}$ and its derivatives, which are fundamental in our proof. Here and subsequently we omit the subscripts $\lambda,\mu$ and use subscripts to denote partial derivatives instead, e.g., $\hat{E}_{\alpha}=\nabla^{\alpha}\hat{E}_{\lambda,\mu}$.
Lemma 2.2.
Suppose $E:\mathbb{Z}^{d}\to\mathbb{R}$ is $\mathbb{Z}^{d}$-symmetric, has vanishing zeroth and second moments as in (2.3), and satisfies $\lvert E(x)\rvert\leq K\lvert x\rvert^{-(d+2+\rho)}$ for some $K,\rho>0$. Choose $\sigma\in(0,\rho)$ such that $\sigma\leq 2$ and let $\alpha$ be a multi-index with $\lvert\alpha\rvert<2+\sigma$. Then there is a constant $c=c(\sigma,\rho,d)$ such that
$\lvert\hat{E}_{\alpha}(k)\rvert\leq cK\cdot\lvert k\rvert^{2+\sigma-\lvert\alpha\rvert}.$ (2.9)
Proof.
Let $g_{x}(k)=\cos(k\cdot x)-1+\frac{(k\cdot x)^{2}}{2!}$. Since $E$ is even, $\sum_{x\in\mathbb{Z}^{d}}x_{i}x_{j}E(x)=0$ for $i\neq j$. By (2.3) and by symmetry, $\sum_{x\in\mathbb{Z}^{d}}x_{i}^{2}E(x)=0$ for all $i$. Therefore,
$\hat{E}(k)=\sum_{x\in\mathbb{Z}^{d}}E(x)\cos(k\cdot x)=\sum_{x\in\mathbb{Z}^{d}}E(x)g_{x}(k).$ (2.10)
By Taylor expansion near $k=0$ and $\sigma\leq 2$, together with elementary properties of sine and cosine, we have $\lvert\nabla^{\alpha}g_{x}(k)\rvert\leq c_{{0}}\lvert k\rvert^{2+\sigma-\lvert\alpha\rvert}\lvert x\rvert^{2+\sigma}$, so
$\lvert\hat{E}_{\alpha}(k)\rvert$ $\leq\sum_{x\in\mathbb{Z}^{d}}\lvert E(x)\rvert\lvert\nabla^{\alpha}g_{x}(k)\rvert\leq c_{{0}}K\cdot\lvert k\rvert^{2+\sigma-\lvert\alpha\rvert}\sum_{x\neq 0}\frac{\lvert x\rvert^{2+\sigma}}{\lvert x\rvert^{d+2+\rho}}.$ (2.11)
The sum on the right-hand side converges since $\sigma<\rho$, and the proof is complete. ∎
For a function $g:\mathbb{T}^{d}\to\mathbb{C}$ and $p\in[1,\infty)$, we denote $L^{p}(\mathbb{T}^{d})$ norms with a subscript $p$:
$\lVert g\rVert_{p}^{p}=\int_{\mathbb{T}^{d}}|g(k)|^{p}\frac{dk}{(2\pi)^{d}},$ (2.12)
and we write $\|g\|_{\infty}$ for the supremum norm.
We will use the following general fact which relates the smoothness of a function on $\mathbb{T}^{d}$ to the decay of its Fourier coefficients. It involves the notion of weak derivative, which is important for functions $h$ which do not decay rapidly enough for $\nabla^{\alpha}\hat{h}$ to exist classically. The weak derivative is defined via integration by parts, and basic facts concerning weak derivatives are recalled in Appendix A. It is by using weak derivatives that we overcome the obstacles to naive application of the ideas in [33] to percolation and LTLA.
Lemma 2.3.
Let $a,d>0$ be positive integers and let $h:\mathbb{Z}^{d}\to\mathbb{R}$. There is a constant $c_{d,a}$, depending only on the dimension $d$ and the maximal order $a$ of differentiation, such that if the weak derivative $\hat{h}_{\alpha}$ is in $L^{1}(\mathbb{T}^{d})$ for all multi-indices $\alpha$ with $|\alpha|\leq a$ then
$|h(x)|\leq c_{d,a}\frac{1}{|\!|\!|x|\!|\!|^{a}}\max_{|\alpha|\in\{0,a\}}\|\hat{h}_{\alpha}\|_{1}.$ (2.13)
Moreover, $|x|^{a}h(x)\to 0$ as $|x|\to\infty$.
Proof.
The bound (2.13) follows from [14, Corollary 3.3.10]. Although stated for classical derivatives, the proof of [14, Corollary 3.3.10] applies for weak derivatives since it only uses integration by parts. Also, it follows from the Riemann–Lebesgue Lemma [14, Proposition 3.3.1] and the integrability of $\hat{h}_{\alpha}$ that the inverse Fourier transform of $\hat{h}_{\alpha}$, which is a constant multiple of $x^{\alpha}h(x)$, vanishes at infinity. Since this holds for any $\lvert\alpha\rvert\leq a$, we obtain $\lvert x\rvert^{a}h(x)\to 0$ as $\lvert x\rvert\to\infty$. ∎
Proposition 2.4.
Let $F$ obey Assumption 1.1. Then the function $\hat{f}=\hat{C}\hat{E}\hat{G}$ is $n_{d}$ times weakly differentiable, and for any multi-index $\alpha$ with $\lvert\alpha\rvert\leq n_{d}$,
$\hat{f}_{\alpha}\in L^{r}(\mathbb{T}^{d})\qquad(r^{-1}>\frac{\lvert\alpha\rvert+2-\rho\wedge 2}{d}).$ (2.14)
Proof of Theorem 2.1 assuming Proposition 2.4.
By Proposition 2.4, $\hat{f}$ is $n_{d}$ times weakly differentiable, so $\nabla^{\alpha}\hat{f}\in L^{1}(\mathbb{T}^{d})$ for all multi-indices $\alpha$ with $\lvert\alpha\rvert\leq n_{d}$. It follows from Lemma 2.3 that $|x|^{n_{d}}f(x)\to 0$ as $\lvert x\rvert\to\infty$. The equation $G=\lambda C_{\mu}+f$ then gives
$G(x)=\lambda C_{{\mu}}(x)+o\left(\frac{1}{\lvert x\rvert^{n_{d}}}\right)\quad\text{as $\lvert x\rvert\to\infty$},$ (2.15)
which completes the proof. ∎
Recall that $\hat{A}=1-\mu\hat{D}$. To prove Proposition 2.4, we must study derivatives of $\hat{f}=\hat{C}\hat{E}\hat{G}=\hat{E}/(\hat{A}\hat{F})$. For this, we would like to use the product and quotient rules of differentiation to conclude that $\hat{f}_{\alpha}$ is given by a linear combination of terms of the form
$\frac{\prod_{n=1}^{i}\hat{A}_{\delta_{n}}}{\hat{A}^{1+i}}\hat{E}_{\alpha_{2}}\frac{\prod_{m=1}^{j}\hat{F}_{\gamma_{m}}}{\hat{F}^{1+j}}=\left(\prod_{n=1}^{i}\frac{\hat{A}_{\delta_{n}}}{\hat{A}}\right)\left(\frac{\hat{E}_{\alpha_{2}}}{\hat{A}\hat{F}}\right)\left(\prod_{m=1}^{j}\frac{\hat{F}_{\gamma_{m}}}{\hat{F}}\right),$ (2.16)
where $\alpha=\alpha_{1}+\alpha_{2}+\alpha_{3}$, $0\leq i\leq\lvert\alpha_{1}\rvert$, $0\leq j\leq\lvert\alpha_{3}\rvert$, $\sum_{n=1}^{i}\delta_{n}=\alpha_{1}$, and $\sum_{m=1}^{j}\gamma_{m}=\alpha_{3}$. However, the use of calculus rules to compute $\hat{f}_{\alpha}$ requires justification. Weak derivatives satisfy a version of product and quotients rules stated in Appendix A. These imply that if a naive application of the product and quotient rules produces a product which a posteriori can be seen to be integrable via Hölder’s inequality, then the weak derivative indeed exists and is given by the usual product and quotient rules. The following lemma is what allows us to use Hölder’s inequality in this manner.
Lemma 2.5.
Let $F$ obey Assumption 1.1. Let $\gamma$ be a multi-index with $\lvert\gamma\rvert<(d-2+\rho\wedge 2)\wedge(\tfrac{1}{2}d+2+\rho)$, and choose $\sigma\in(0,\rho)$ such that $\sigma\leq 2$. Choose $q_{1},q_{2}$ satisfying
$\frac{\lvert\gamma\rvert}{d}<q_{1}^{-1},\qquad\frac{2-\sigma+\lvert\gamma\rvert}{d}<q_{2}^{-1}.$ (2.17)
Then $\hat{F},\hat{A},\hat{E}$ are $\gamma$-times weakly differentiable, and
$\frac{\hat{A}_{\gamma}}{\hat{A}},\;\frac{\hat{F}_{\gamma}}{\hat{F}}\in L^{q_{1}}(\mathbb{T}^{d}),\qquad\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}}\in L^{q_{2}}(\mathbb{T}^{d}).$ (2.18)
Proof of Proposition 2.4 assuming Lemma 2.5.
Let $\lvert\alpha\rvert\leq n_{d}$, $\rho_{2}=\rho\wedge 2$, and pick some $\sigma\in(0,\rho_{2})$. We use the quotient and product rules (Lemmas A.2 and A.3) to calculate $\hat{f}_{\alpha}$, and we use Lemma 2.5 to verify the hypotheses of the rules. We need to verify all terms of the form (2.16) are integrable to conclude that $\hat{f}_{\alpha}$ exists and is given by a linear combination of (2.16). By Hölder’s inequality and Lemma 2.5, the product (2.16) belongs to $L^{r}(\mathbb{T}^{d})$ as long as
$\frac{1}{r}>\frac{\sum_{n=1}^{i}\lvert\delta_{n}\rvert}{d}+\frac{2-\sigma+\lvert\alpha_{2}\rvert}{d}+\frac{\sum_{m=1}^{j}\lvert\gamma_{m}\rvert}{d}=\frac{\lvert\alpha\rvert+2-\sigma}{d}.$ (2.19)
[…]
To complete the proof of Theorem 2.1, it remains to prove Lemma 2.5. The proof uses the following elementary facts.
Lemma 2.6.
Let $h:\mathbb{Z}^{d}\to\mathbb{R}$ obey $|h(x)|\leq K|\!|\!|x|\!|\!|^{-b}$ for some $K,b>0$.
  1. (i)
    If $b>d$ then $h\in\ell^{1}(\mathbb{Z}^{d})$, $\hat{h}\in L^{\infty}(\mathbb{T}^{d})$, and $\lVert\hat{h}\rVert_{\infty}\leq c_{d,b}K$.
  2. (ii)
    If $b\leq d$ then $h\in\ell^{p}(\mathbb{Z}^{d})$ for $p>d/b$. If also $\frac{d}{2}<b\leq d$ then $\hat{h}\in L^{q}(\mathbb{T}^{d})$ and $\lVert\hat{h}\rVert_{q}\leq c_{d,b,q}K$ for all $1\leq q<d/(d-b)$.
Proof.
(i) This follows from the summability of $|x|^{-b}$.
(ii) The fact that $h\in\ell^{p}(\mathbb{Z}^{d})$ for $p>d/b$ follows from the summability of $|x|^{-pb}$ for $pb>d$. The rest follows from the fact that the Fourier transform is bounded as a map from $\ell^{p}(\mathbb{Z}^{d})\to L^{q}(\mathbb{T}^{d})$ when $p\in[1,2]$ and $p^{-1}+q^{-1}=1$, which requires $\frac{d}{b}<2$ and $2\leq q<\frac{d}{d-b}$. The extension to $1\leq q<\frac{d}{d-b}$ follows from the monotonicity of the $L^{q}(\mathbb{T}^{d})$ norm in $q$. ∎
For example, if $h(x)$ decays like $|x|^{-(d+2+\rho)}$ then $x^{\alpha}h(x)$ decays like $|x|^{-(d+2+\rho-|\alpha|)}$. If $\lvert\alpha\rvert<2+\rho$ then Lemma 2.6(i) implies $\nabla^{\alpha}\hat{h}\in L^{\infty}(\mathbb{T}^{d})$. If $\lvert\alpha\rvert\geq 2+\rho$, then $|x|^{-(d+2+\rho-|\alpha|)}$ is in $\ell^{p}(\mathbb{Z}^{d})$ for $p>\frac{d}{d+2+\rho-\lvert\alpha\rvert}$, and we can take $p=2$ provided
$\lvert\alpha\rvert<\tfrac{1}{2}d+2+\rho.$ (2.20)
In this case, Lemma A.4 implies that $\hat{h}$ is $\alpha$-times weakly differentiable, and Lemma 2.6(ii) implies that $\nabla^{\alpha}\hat{h}\in L^{q}(\mathbb{T}^{d})$ for $1\leq q<d/(\lvert\alpha\rvert-2-\rho)$. When $|\alpha|=d-2$, (2.20) is the origin of our restriction $\rho>\frac{d-8}{2}$, which we believe to be artificial in the deconvolution theorem.
2.2.2 Proof of Lemma 2.5
Bound on $\hat{A}_{\gamma}/\hat{A}$. By definition, $\hat{A}=1-\mu\hat{D}$. Let $\gamma$ be any multi-index. There is nothing to prove for $\lvert\gamma\rvert=0$ since the ratio is then $1$. If $\lvert\gamma\rvert=1$, by Taylor’s theorem and symmetry, we have $\lvert\hat{A}_{\gamma}(k)\rvert\lesssim\lvert k\rvert$. If $\lvert\gamma\rvert\geq 2$, Taylor’s theorem gives $\lvert\hat{A}_{\gamma}(k)\rvert\lesssim 1$ instead. It then follows from the infrared bound (1.11) that
$\Big{\lvert}\frac{\hat{A}_{\gamma}}{\hat{A}}(k)\Big{\rvert}\lesssim\lvert k\rvert^{-(\lvert\gamma\rvert\wedge 2)}\in L^{q}(\mathbb{T}^{d})\qquad(q^{-1}>\frac{\lvert\gamma\rvert\wedge 2}{d}),$ (2.21)
which is stronger than the desired (2.18).
Bound on $\hat{F}_{\gamma}/\hat{F}$. Let $\lvert\gamma\rvert<\tfrac{1}{2}d+2+\rho$. There is again nothing to prove for $\lvert\gamma\rvert=0$. If $\lvert\gamma\rvert=1$, by Taylor’s theorem and symmetry, and since $\sum_{x\in\mathbb{Z}^{d}}|x|^{2}|F(x)|$ is finite, we have $\lvert\hat{F}_{\gamma}(k)\rvert\lesssim\lvert k\rvert$. It then follows from the infrared bound that $\lvert\hat{F}_{\gamma}(k)/\hat{F}(k)\rvert\lesssim\lvert k\rvert^{-1}\in L^{q}$ for $q^{-1}>1/d$.
For $\lvert\gamma\rvert\geq 2$, by the decay hypothesis on $F(x)$, $\hat{F}_{\gamma}\in L^{\infty}$ exists classically if $|\gamma|<2+\rho$, and $\hat{F}_{\gamma}$ exists as a weak derivative (see the discussion around (2.20)) if $2+\rho\leq|\gamma|<\frac{1}{2}d+2+\rho$. In the second case, we also have $\hat{F}_{\gamma}\in L^{r}$ whenever $r^{-1}>(\lvert\gamma\rvert-2-\rho)/d$ by Lemma 2.6. Since $\rho>0$, we get
$\hat{F}_{\gamma}\in L^{d/(\lvert\gamma\rvert-2)}(\mathbb{T}^{d})\qquad(2\leq\lvert\gamma\rvert<\tfrac{1}{2}d+2+\rho).$ (2.22)
Since $\lvert\hat{F}^{-1}(k)\rvert\lesssim\lvert k\rvert^{-2}\in L^{p}$ for all $p^{-1}>2/d$ by the infrared bound, it follows from Hölder’s inequality that $\hat{F}_{\gamma}/\hat{F}\in L^{q}$ for all $q^{-1}>(\lvert\gamma\rvert-2+2)/d$, as desired.
Bound on $\hat{E}_{\gamma}/\hat{A}\hat{F}$. Let $\lvert\gamma\rvert<\tfrac{1}{2}d+2+\rho$. We divide into two cases, now according to how $\lvert\gamma\rvert$ compares with $2+\sigma$. We use the fact that $E=A-\lambda F$ has the same $|x|^{-(d+2+\rho)}$ decay as $F$.
If $\lvert\gamma\rvert<2+\sigma$, Lemma 2.2 and the infrared bounds give $\lvert\hat{E}_{\gamma}/(\hat{A}\hat{F})(k)\rvert\lesssim\lvert k\rvert^{2+\sigma-\lvert\gamma\rvert-4}=\lvert k\rvert^{-(2-\sigma+\lvert\gamma\rvert)}$, which is in $L^{q}$ for $q^{-1}>(2-\sigma+\lvert\gamma\rvert)/d$.
If $2+\sigma\leq\lvert\gamma\rvert<\frac{1}{2}d+2+\rho$, as in the proof for $\hat{F}_{\gamma}/\hat{F}$, from the decay of $E(x)$ we get $\hat{E}_{\gamma}\in L^{\infty}$ when $2+\sigma\leq\lvert\gamma\rvert<2+\rho$, and $\hat{E}_{\gamma}\in L^{r}$ for $r^{-1}>(\lvert\gamma\rvert-2-\rho)/d$ when $2+\rho\leq\lvert\gamma\rvert<\frac{1}{2}d+2+\rho$. Since $\sigma<\rho$, in both subcases we have
$\hat{E}_{\gamma}\in L^{r}(\mathbb{T}^{d})\qquad(r^{-1}>\frac{\lvert\gamma\rvert-2-\sigma}{d}).$ (2.23)
The desired result then follows from Hölder’s inequality and the infrared bound $\lvert(\hat{A}\hat{F})^{-1}(k)\rvert\lesssim\lvert k\rvert^{-4}\in L^{p}$ for all $p^{-1}>4/d$. ∎
For later use, we note that the same reasoning gives, for $\sigma\leq\lvert\gamma\rvert<\frac{1}{2}d+2+\rho$,
$\frac{\hat{E}_{\gamma}}{\hat{A}},\;\frac{\hat{E}_{\gamma}}{\hat{F}}\in L^{q}(\mathbb{T}^{d})\qquad(q^{-1}>\frac{\lvert\gamma\rvert-\sigma}{d}).$ (2.24)
2.3 Improved error estimate for Theorem 1.2
We now improve the error estimates of Section 2.2 using an integral representation of a fractional power of $t\geq 0$, namely that for $\delta\in(0,1)$,
[…]
This representation is similar to but not the same as that of [9, Section 4.2]. In the Fourier space, the fractional power has the effect of a fractional derivative.
2.3.1 Fractional derivatives
Given a function $\hat{g}:\mathbb{T}^{d}\to\mathbb{C}$, we regard it as a periodic function on $\mathbb{R}^{d}$ and define
$(U_{u}\hat{g})(k)=\hat{g}(k+\tilde{u})-\hat{g}(k-\tilde{u})$ (2.26)
with $\tilde{u}=(u,0,\dots,0)\in\mathbb{R}^{d}$.
Lemma 2.7.
[…]
Then the inverse Fourier transform of $\hat{w}(k)$ is $w(x)=({\rm sgn}\,x_{1})|x_{1}|^{\delta}g(x)$.
Proof.
By hypothesis, $\hat{w}\in L^{1}(\mathbb{T}^{d})$, so its inverse Fourier transform is
$w(x)$ $=\int_{\mathbb{T}^{d}}\hat{w}(k)e^{-ik\cdot x}\frac{dk}{(2\pi)^{d}}=\frac{1}{2ic_{\delta}}\int_{0}^{\infty}\frac{du}{u^{1+\delta}}\int_{\mathbb{T}^{d}}\frac{dk}{(2\pi)^{d}}(U_{u}\hat{g})(k)e^{-ik\cdot x},$ (2.28)
where the application of Fubini’s Theorem is justified by hypothesis. A change of variables to evaluate the $k$-integral gives
$w(x)$ $=\frac{1}{c_{\delta}}g(x)\int_{0}^{\infty}\frac{du}{u^{1+\delta}}\sin(ux_{1}),$ (2.29)
and (2.25) then gives $w(x)=({\rm sgn}\,x_{1})|x_{1}|^{\delta}g(x)$. ∎
The following lemma improves Lemma 2.3 by a fractional power of $|\!|\!|x|\!|\!|$.
Lemma 2.8.
Let $d>0$ and let $h:\mathbb{Z}^{d}\to\mathbb{R}$ be $\mathbb{Z}^{d}$-symmetric. Suppose that $a>0$ is not an integer, and that the weak derivative $\hat{h}_{\alpha}$ is in $L^{1}(\mathbb{T}^{d})$ for all multi-indices $\alpha$ with $|\alpha|=\lfloor a\rfloor$. Suppose also that, for some $\eta\in(a-\lfloor a\rfloor,1)$ and for all $\lvert\alpha\rvert=\lfloor a\rfloor$,
$\lVert U_{u}\hat{h}_{\alpha}\rVert_{1}\leq u^{\eta}K_{\alpha}\qquad(0\leq u\leq 1).$ (2.30)
Then there is a constant $c_{d,a,\eta}$ depending only on $d$, $a$, and $\eta$ such that
$\lvert h(x)\rvert\leq c_{d,a,\eta}\frac{1}{|\!|\!|x|\!|\!|^{a}}\Big{(}\|\hat{h}\|_{1}+\max_{\lvert\alpha\rvert=\lfloor a\rfloor}(K_{\alpha}+\lVert\hat{h}_{\alpha}\rVert_{1})\Big{)}.$ (2.31)
Proof.
We write $\delta=a-\lfloor a\rfloor\in(0,\eta)$. For $x=0$, there is nothing to prove since the bound $\lvert h(0)\rvert\leq\lVert\hat{h}\rVert_{1}$ implies (2.31). Let $x\neq 0$ and $|\alpha|=\lfloor a\rfloor$. We first use the hypothesis to observe that
[…]
$|x_{1}|^{\delta}|x^{\alpha}h(x)|=|w(x)|\leq\|\hat{w}\|_{1}\lesssim K_{\alpha}+\|h_{\alpha}\|_{1}.$ (2.33)
By symmetry, the above also holds with $x_{1}$ replaced by $x_{j}$ for any $j$. Since $a=|\alpha|+\delta$, we have $|x|^{a}|h(x)|\lesssim\max_{|\alpha|=\lfloor a\rfloor}(K_{\alpha}+\|h_{\alpha}\|_{1})$. Together with the bound for $x=0$, this completes the proof. ∎
When it is possible to take a full derivative of $\hat{h}_{\alpha}$, we can verify (2.30) using the next lemma, which can be considered as an estimate on a fractional difference quotient. For $\mathbb{R}^{d}$, the case $\eta=1$ can be found, e.g., in [10, Sec. 5.8.2, Theorem 3].
Lemma 2.9.
Let $g:\mathbb{T}^{d}\to\mathbb{C}$ be weakly differentiable. Fix $1\leq p<d$. Assume $\nabla^{e_{j}}g\in L^{p}(\mathbb{T}^{d})$ for all $j$. Let $0\leq\eta\leq 1$ and define $p_{\eta}$ by $\frac{1}{p_{\eta}}=\frac{1}{p}-\frac{1-\eta}{d}$. Then
$\lVert U_{u}g\rVert_{p_{\eta}}\lesssim u^{\eta}\lVert g\rVert_{W^{1,p}},$ (2.34)
where $\lVert g\rVert_{W^{1,p}}=(\lVert g\rVert_{p}^{p}+\sum_{j=1}^{d}\lVert\nabla_{j}g\rVert_{p}^{p})^{1/p}$.
Proof.
Define $p^{*}$ by $\frac{1}{p^{*}}=\frac{1}{p}-\frac{1}{d}$, so that $\frac{1}{p_{\eta}}=\frac{\eta}{p}+\frac{1-\eta}{p^{*}}$. Since $\lVert U_{u}g\rVert_{p_{\eta}}\leq\lVert U_{u}g\rVert_{p}^{\eta}\lVert U_{u}g\rVert_{p^{*}}^{1-\eta}$ by the log-convexity of $L^{p}$ norms, it suffices to prove the cases $\eta=0,1$. The case $\eta=0$ follows from the triangle inequality and the torus Sobolev inequality [3, Corollary 1.2]:
$\lVert U_{u}g\rVert_{p^{*}}\leq 2\lVert g\rVert_{p^{*}}\lesssim\lVert g\rVert_{W^{1,p}}.$ (2.35)
For $\eta=1$, since $C^{1}$ is dense in $W^{1,p}$ it suffices to consider $g\in C^{1}$, for which the Fundamental Theorem of Calculus gives
[…]
Then by Minkowski’s integral inequality,
$\lVert U_{u}g\rVert_{p}\leq u\int_{-1}^{1}\lVert\nabla^{e_{1}}g(\cdot+t\tilde{u})\rVert_{p}dt=2u\lVert\nabla^{e_{1}}g\rVert_{p}\leq 2u\lVert g\rVert_{W^{1,p}}.$ (2.37)
This completes the proof. ∎
When it is not possible to take a full derivative as required by Lemma 2.9, we use the following elementary lemma.
Lemma 2.10.
Let $h:\mathbb{Z}^{d}\to\mathbb{R}$ obey $|h(x)|\leq K|\!|\!|x|\!|\!|^{-b}$ for some $K,b>0$, and let $0\leq\eta\leq 1$.
  1. (i)
    If $b>d+\eta$ then $\hat{h}\in L^{\infty}(\mathbb{T}^{d})$ and $\lVert U_{u}\hat{h}\rVert_{\infty}\leq c_{b,\eta}u^{\eta}K$.
  2. (ii)
    If $\frac{1}{2}d+\eta<b\leq d+\eta$ then $\hat{h}\in L^{q}(\mathbb{T}^{d})$ and $\lVert U_{u}\hat{h}\rVert_{q}\leq c_{d,b,q}u^{\eta}K$ for all $1\leq q<d/(d-b+\eta)$.
Proof.
(i) The claim that $\hat{h}\in L^{\infty}(\mathbb{T}^{d})$ follows from the summability of $|x|^{-b}$. For the claim on $U_{u}\hat{h}$, we use $\lvert\sin(ux_{1})\rvert\lesssim\lvert ux_{1}\rvert^{\eta}\leq u^{\eta}\lvert x\rvert^{\eta}$ and the summability of $\lvert x\rvert^{\eta-b}$ to see that
$\lVert U_{u}\hat{h}\rVert_{\infty}\leq\lVert 2i\sin(ux_{1})h(x)\rVert_{1}\lesssim u^{\eta}\lVert\lvert x\rvert^{\eta}h(x)\rVert_{1}\lesssim u^{\eta}K.$ (2.38)
(ii) Note that $\lvert x\rvert^{\eta}\lvert h(x)\rvert\leq K\lvert x\rvert^{\eta-b}$ is in $\ell^{p}(\mathbb{Z}^{d})$ for $p>d/(b-\eta)$. The condition $b>\frac{1}{2}d+\eta$ ensures that we can take $p$ less than $2$. Using the same bound on $\sin(ux_{1})$ as in part (i) and boundness of the Fourier transform as a map from $\ell^{p}(\mathbb{Z}^{d})\to L^{q}(\mathbb{T}^{d})$, with $p\in[1,2]$ and $p^{-1}+q^{-1}=1$, we get
$\lVert U_{u}\hat{h}\rVert_{q}\leq\lVert 2i\sin(ux_{1})h(x)\rVert_{p}\lesssim u^{\eta}\lVert\lvert x\rvert^{\eta}h(x)\rVert_{p}\lesssim u^{\eta}K$ (2.39)
for $2\leq q<\frac{d}{d-b+\eta}$. The extension to $1\leq q<\frac{d}{d-b+\eta}$ follows from the monotonicity of the $L^{q}(\mathbb{T}^{d})$ norm in $q$. The claim on $\hat{h}$ is similar and does not use $\sin(ux_{1})$. ∎
2.3.2 Proof of Theorem 1.2
With Lemma 2.8, we reduce the proof of Theorem 1.2 to the following proposition.
Proposition 2.11.
Let $\lvert\alpha\rvert\leq n_{d}$, $0\leq u\leq 1$, and suppose $\eta\in(0,1)$ satisfies
$n_{d}+\eta<(d-2+\rho)\wedge(\tfrac{1}{2}d+2+\rho).$ (2.40)
If $F$ obeys Assumption 1.1, then
$\lVert U_{u}\hat{f}_{\alpha}\rVert_{1}\lesssim u^{\eta}.$ (2.41)
Proof of Theorem 1.2 assuming Proposition 2.11.
By definition, $s\in[0,\rho\wedge 2\wedge(\rho-\frac{d-8}{2}))$. It suffices to prove the theorem for values of $s$ close to the upper limit of this interval, so we assume $s$ is not an integer. Let $\delta=s-\lfloor s\rfloor\in(0,1)$. Then
$d-2+s=n_{d}+\delta<(d-2+\rho\wedge 2)\wedge(\tfrac{1}{2}d+2+\rho).$ (2.42)
Since we have a strict inequality, we can pick an $\eta\in(\delta,1)$ satisfying (2.40). It then follows from Lemma 2.8 (with $a=n_{d}+\delta$), Proposition 2.4, and Proposition 2.11 that $f(x)=O(|\!|\!|x|\!|\!|^{-(d-2+s)})$, and the desired result follows from $G=\lambda C_{\mu}+f$. ∎
To prove Proposition 2.11, as in (2.16), we write $\hat{f}_{\alpha}$ as a linear combination of terms of the form
$\left(\prod_{n=1}^{i}\frac{\hat{A}_{\delta_{n}}}{\hat{A}}\right)\left(\frac{\hat{E}_{\alpha_{2}}}{\hat{A}\hat{F}}\right)\left(\prod_{m=1}^{j}\frac{\hat{F}_{\gamma_{m}}}{\hat{F}}\right).$ (2.43)
To calculate $U_{u}\hat{f}_{\alpha}$, we use the linearity of $U_{u}$ and apply the following elementary formula for the difference of two products. Let $X_{k,l}=\prod_{i=k}^{l}a_{i}$ and $Y_{k,l}=\prod_{i=k}^{l}b_{i}$. Then, with empty products equal to $1$,
$X_{1,n}-Y_{1,n}=\sum_{j=1}^{n}(a_{j}-b_{j})X_{1,j-1}Y_{j+1,n}.$ (2.44)
We apply this to (2.43), so that $U_{u}$ acts on one factor at a time. We will prove the following lemma.
Lemma 2.12.
Let $\gamma$ be a multi-index, $0\leq\eta\leq 1$, with $\lvert\gamma\rvert+\eta<(d-2+\rho\wedge 2)\wedge(\tfrac{1}{2}d+2+\rho)$. Choose $\sigma\in(0,\rho)$ such that $\sigma\leq 2$, and choose $q_{1},q_{2}$ satisfying
$\frac{\lvert\gamma\rvert+\eta}{d}<q_{1}^{-1},\qquad\frac{2-\sigma+\lvert\gamma\rvert+\eta}{d}<q_{2}^{-1}.$ (2.45)
Then, for $F$ obeying Assumption 1.1 and for $0\leq u\leq 1$,
$\Big{\lVert}U_{u}(\frac{\hat{A}_{\gamma}}{\hat{A}})\Big{\rVert}_{q_{1}},\;\Big{\lVert}U_{u}(\frac{\hat{F}_{\gamma}}{\hat{F}})\Big{\rVert}_{q_{1}},\;\Big{\lVert}U_{u}(\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}})\Big{\rVert}_{q_{2}}\lesssim u^{\eta}.$ (2.46)
Proof of Proposition 2.11 assuming Lemma 2.12.
By hypothesis, $n_{d}+\eta<(d-2+\rho)\wedge(\tfrac{1}{2}d+2+\rho)$. Since $n_{d}+\eta<d-1+1=d$, we can pick $\sigma$ slightly smaller than $\rho\wedge 2$ so that we still have $n_{d}+\eta<d-2+\sigma$.
To bound $U_{u}\hat{f}_{\alpha}$, it suffices to bound the $L^{1}$ norm of $U_{u}$ applied to (2.43). We use (2.44) and bound the $L^{1}$ norm of, e.g.,
$\left(\prod_{n=1}^{i}\frac{\hat{A}_{\delta_{n}}}{\hat{A}}(k+\tilde{u})\right)U_{u}(\frac{\hat{E}_{\alpha_{2}}}{\hat{A}\hat{F}})(k)\left(\prod_{m=1}^{j}\frac{\hat{F}_{\gamma_{m}}}{\hat{F}}(k-\tilde{u})\right).$ (2.47)
Recall from Lemma 2.5 that, under our current hypotheses,
$\frac{\hat{F}_{\gamma}}{\hat{F}},\frac{\hat{A}_{\gamma}}{\hat{A}}$ $\in L^{q}\quad(q^{-1}>\frac{\lvert\gamma\rvert}{d}),\qquad\frac{\hat{E}_{\alpha}}{\hat{A}\hat{F}}\in L^{q}\quad(q^{-1}>\frac{2-\sigma+\lvert\alpha\rvert}{d}).$ (2.48)
By Hölder’s inequality, the $L^{r}$ norm of (2.47) is bounded by
$\left(\prod_{n=1}^{i}\Big{\lVert}\frac{\hat{A}_{\delta_{n}}}{\hat{A}}\Big{\rVert}_{r_{n}}\right)\Big{\lVert}U_{u}(\frac{\hat{E}_{\alpha_{2}}}{\hat{A}\hat{F}})\Big{\rVert}_{q}\left(\prod_{m=1}^{j}\Big{\lVert}\frac{\hat{F}_{\gamma_{m}}}{\hat{F}}\Big{\rVert}_{q_{m}}\right),$ (2.49)
where $r$ can be any number that satisfies
$\frac{1}{r}$ $=\left(\sum_{n=1}^{i}\frac{1}{r_{n}}\right)+\frac{1}{q}+\left(\sum_{m=1}^{j}\frac{1}{q_{m}}\right)$
$>\frac{\sum_{n=1}^{i}\lvert\delta_{n}\rvert}{d}+\frac{2-\sigma+\lvert\alpha_{2}\rvert+\eta}{d}+\frac{\sum_{m=1}^{j}\lvert\gamma_{m}\rvert}{d}=\frac{\lvert\alpha\rvert+\eta+2-\sigma}{d}.$ (2.50)
In the above, $r_{n}$ and $q_{m}$ are dictated by Lemma 2.5, and $q$ is dictated by Lemma 2.13. Since $\lvert\alpha\rvert+\eta\leq n_{d}+\eta<d-2+\sigma$, we can take $r=1$. The norm contains a factor $u^{\eta}$ from $\lVert U_{u}(\frac{\hat{E}_{\alpha_{2}}}{\hat{A}\hat{F}})\rVert_{q}$. Similar terms with $U_{u}$ instead applied to one of the $\hat{F}_{\gamma}/\hat{F}$ or $\hat{A}_{\gamma}/\hat{A}$ factors are treated in the same way, using Lemma 2.12. ∎
Before we prove Lemma 2.12, we give a further consequence of the good moment behaviour of $E$ in (2.3). The next lemma is similar to Lemma 2.2 and bounds $U_{u}\hat{E}$ and its derivatives. Instead of the good powers of $\lvert k\rvert$ in (2.9), we now have good powers of $\lvert k\pm\tilde{u}\rvert$.
Lemma 2.13.
Suppose $E:\mathbb{Z}^{d}\to\mathbb{R}$ is $\mathbb{Z}^{d}$-symmetric, has vanishing zeroth and second moments as in (2.3), and satisfies $\lvert E(x)\rvert\leq K\lvert x\rvert^{-(d+2+\rho)}$ for some $K,\rho>0$. Choose $\sigma\in(0,\rho)$ with $\sigma\leq 2$, choose $0\leq\eta\leq 1$, and let $\alpha$ be a multi-index with $\lvert\alpha\rvert+\eta<2+\sigma$. Then there is a constant $c=c(\sigma,\rho,d)$ such that
$\lvert U_{u}\hat{E}_{\alpha}(k)\rvert\leq cK\cdot u^{\eta}(\lvert k+\tilde{u}\rvert^{2+\sigma-\eta-\lvert\alpha\rvert}+\lvert k-\tilde{u}\rvert^{2+\sigma-\eta-\lvert\alpha\rvert}).$ (2.51)
Proof.
We prove that
$\lvert U_{u}\hat{E}_{\alpha}(k)\rvert\leq cK\cdot\begin{cases}u^{1+\sigma}\lvert k\rvert+u\lvert k\rvert^{1+\sigma}&(\lvert\alpha\rvert=0)\\ u^{2+\sigma-\lvert\alpha\rvert}+u\lvert k\rvert^{1+\sigma-\lvert\alpha\rvert}&(1\leq\lvert\alpha\rvert<1+\sigma)\\ u^{\sigma}&(\sigma\leq 1,\;\lvert\alpha\rvert=2)\\ u^{\sigma-1}&(1<\sigma\leq 2,\;\lvert\alpha\rvert=3).\end{cases}$ (2.52)
The claim (2.51) then follows from the inequalities $\lvert k\rvert,u\leq 2^{-1}(\lvert k+\tilde{u}\rvert+\lvert k-\tilde{u}\rvert)$ and $(\lvert a\rvert+\lvert b\rvert)^{p}\leq 2^{p}(\lvert a\rvert^{p}+\lvert b\rvert^{p})$ with $p=2+\sigma-\eta-\lvert\alpha\rvert\geq 0$.
For the first two lines of (2.52), we use the identity
$\frac{1}{2i}U_{u}\hat{E}_{\alpha}(k)=\sum_{x\in\mathbb{Z}^{d}}\big{(}\sin(ux_{1})-ux_{1}\big{)}(ix)^{\alpha}E(x)e^{ik\cdot x}+\sum_{x\in\mathbb{Z}^{d}}ux_{1}(ix)^{\alpha}E(x)e^{ik\cdot x}.$ (2.53)
The second term is exactly $\frac{u}{i}\hat{E}_{\alpha+e_{1}}(k)$ and is bounded in absolute value by a multiple of $Ku\lvert k\rvert^{1+\sigma-\lvert\alpha\rvert}$ by Lemma 2.2. If $\lvert\alpha\rvert=0$, we bound the first term of (2.53) using symmetry. With $\tilde{x}=(x_{2},\dots,x_{d})$ and $\tilde{k}=(k_{2},\dots,k_{d})$,
$\Big{\lvert}\sum_{x\in\mathbb{Z}^{d}}\big{(}\sin(ux_{1})-ux_{1}\big{)}E(x)e^{ik\cdot x}\Big{\rvert}$ $=\Big{\lvert}\sum_{x\in\mathbb{Z}^{d}}\big{(}\sin(ux_{1})-ux_{1}\big{)}E(x)e^{i\tilde{k}\cdot\tilde{x}}\sin(k_{1}x_{1})\Big{\rvert}$
$\leq c\sum_{x\in\mathbb{Z}^{d}}\lvert ux_{1}\rvert^{1+\sigma}\lvert E(x)\rvert\lvert k_{1}x_{1}\rvert\lesssim Ku^{1+\sigma}\lvert k\rvert,$ (2.54)
because $1+\sigma\leq 3$ and $E(x)$ has enough decay. If $\lvert\alpha\rvert\geq 1$, then $2+\sigma-\lvert\alpha\rvert\leq 3$. We instead use
$\Big{\lvert}\sum_{x\in\mathbb{Z}^{d}}\big{(}\sin(ux_{1})-ux_{1}\big{)}(ix)^{\alpha}E(x)e^{ik\cdot x}\Big{\rvert}$ $\leq c\sum_{x\in\mathbb{Z}^{d}}\lvert ux_{1}\rvert^{2+\sigma-\lvert\alpha\rvert}\lvert x\rvert^{\lvert\alpha\rvert}\lvert E(x)\rvert\lesssim Ku^{2+\sigma-\lvert\alpha\rvert}.$ (2.55)
For the last two lines of (2.52), we bound $\sin(ux_{1})$ directly. Since $2+\sigma-\lvert\alpha\rvert\leq 1$, we get
$\lvert U_{u}\hat{E}_{\alpha}(k)\rvert$ $\leq 2c\sum_{x\in\mathbb{Z}^{d}}\lvert ux_{1}\rvert^{2+\sigma-\lvert\alpha\rvert}\lvert x\rvert^{\lvert\alpha\rvert}\lvert E(x)\rvert\lesssim Ku^{2+\sigma-\lvert\alpha\rvert}$ (2.56)
by the decay of $E(x)$. ∎
2.3.3 Proof of Lemma 2.12
It remains to prove Lemma 2.12. We start by proving (2.46) for $\hat{F}$. The proof for $\hat{A}$ is analogous so we omit it.
Bound on $U_{u}(\hat{F}_{\gamma}/\hat{F})$. The statement is trivial for $\lvert\gamma\rvert=0$, as $U_{u}(1)=0$. We treat the cases $\lvert\gamma\rvert=1$ and $\lvert\gamma\rvert\geq 2$ separately.
Case 1: $\lvert\gamma\rvert=1$. Let $i,j\in\{1,\dots,d\}$ and write $\gamma=e_{i}$. We use Lemma 2.9. It follows from the quotient rule, the infrared bound, $\nabla_{ij}\hat{F}\in L^{\infty}$, and $\lvert\nabla_{i}\hat{F}(k)\rvert,\lvert\nabla_{i}\hat{F}(k)\rvert\lesssim\lvert k\rvert$ that
$\Big{\lvert}\nabla_{j}(\frac{\hat{F}_{\gamma}}{\hat{F}})\Big{\rvert}=\Big{\lvert}\frac{\nabla_{ij}\hat{F}}{\hat{F}}-\frac{\nabla_{i}\hat{F}\nabla_{j}\hat{F}}{\hat{F}^{2}}\Big{\rvert}\lesssim\frac{1}{\lvert k\rvert^{2}},$ (2.57)
which is in $L^{p}$ for $p^{-1}>2/d$. Since $d>2$, this implies that $\hat{F}_{\gamma}/\hat{F}$ is weakly differentiable, and Lemma 2.9 gives $\lVert U_{u}(\hat{F}_{\gamma}/\hat{F})\rVert_{q}\lesssim u^{\eta}$ for
$\frac{1}{q}=\frac{1}{p_{\eta}}=\frac{1}{p}-\frac{1-\eta}{d}>\frac{1+\eta}{d},$ (2.58)
which is the desired result.
Case 2: $2\leq\lvert\gamma\rvert<\frac{1}{2}d+2+\rho$. By (2.44),
$U_{u}(\frac{\hat{F}_{\gamma}}{\hat{F}})(k)=(U_{u}\hat{F}_{\gamma})(k)\hat{F}^{-1}(k+\tilde{u})+\hat{F}_{\gamma}(k-\tilde{u})U_{u}\hat{F}^{-1}(k).$ (2.59)
We show each of the two terms on the right-hand side has $L^{q}$ norm at most $u^{\eta}$ when $q^{-1}>(\lvert\gamma\rvert+\eta)/d$.
For the first term of (2.59), since $\lvert\gamma\rvert+\eta<\frac{1}{2}d+2+\rho$, applying Lemma 2.10 to $x^{\gamma}F(x)$, which decays like $\lvert x\rvert^{-(d+2+\rho-\lvert\gamma\rvert)}$, gives
$\lVert U_{u}\hat{F}_{\gamma}\rVert_{\frac{d}{\lvert\gamma\rvert+\eta-2}}\lesssim u^{\eta}$ (2.60)
when $\lvert\gamma\rvert\geq 2$. Since the infrared bound $\lvert\hat{F}^{-1}(k+\tilde{u})\rvert\lesssim\lvert k+\tilde{u}\rvert^{-2}$ implies that $\hat{F}^{-1}(k+\tilde{u})\in L^{r}$ for $r^{-1}>2/d$, Hölder’s inequality gives $\lVert(U_{u}\hat{F}_{\gamma})(k)\hat{F}^{-1}(k+\tilde{u})\rVert_{q}\lesssim u^{\eta}$ for $q^{-1}>(\lvert\gamma\rvert+\eta)/d$, as desired.
For the second term of (2.59), we recall $\hat{F}_{\gamma}\in L^{d/(\lvert\gamma\rvert-2)}$ from (2.22). To calculate $U_{u}\hat{F}^{-1}$, we use the symmetry of $F(x)$ to write
$U_{u}\hat{F}(k)$ $=2i\sum_{x\in\mathbb{Z}^{d}}\sin(ux_{1})F(x)e^{ik\cdot x}=2i^{2}\sum_{x\in\mathbb{Z}^{d}}\sin(ux_{1})F(x)e^{i\tilde{k}\cdot\tilde{x}}\sin(k_{1}x_{1}),$ (2.61)
where $\tilde{x}=(x_{2},\dots,x_{d})$ and $\tilde{k}=(k_{2},\dots,k_{d})$. Since $\lvert\sin t\rvert\leq\lvert t\rvert$,
$\lvert U_{u}\hat{F}(k)\rvert$ $\leq 2\sum_{x\in\mathbb{Z}^{d}}u\lvert x_{1}\rvert\lvert F(x)\rvert\lvert k_{1}x_{1}\rvert\lesssim u\lvert k\rvert.$ (2.62)
By the inequalities $\lvert k\rvert,u\leq 2^{-1}(\lvert k+\tilde{u}\rvert+\lvert k-\tilde{u}\rvert)$ and $(\lvert a\rvert+\lvert b\rvert)^{p}\leq 2^{p}(\lvert a\rvert^{p}+\lvert b\rvert^{p})$ with $p=2-\eta\geq 0$,
$\lvert U_{u}\hat{F}(k)\rvert\lesssim u\lvert k\rvert\leq u^{\eta}(\lvert k+\tilde{u}\rvert^{2-\eta}+\lvert k-\tilde{u}\rvert^{2-\eta}),$ (2.63)
and hence
$\lvert(U_{u}\hat{F}^{-1})(k)\rvert$ $=\left\lvert\frac{-U_{u}\hat{F}(k)}{\hat{F}(k+\tilde{u})\hat{F}(k-\tilde{u})}\right\rvert\lesssim u^{\eta}\left(\frac{1}{\lvert k+\tilde{u}\rvert^{\eta}\lvert k-\tilde{u}\rvert^{2}}+\frac{1}{\lvert k+\tilde{u}\rvert^{2}\lvert k-\tilde{u}\rvert^{\eta}}\right).$ (2.64)
We use Hölder’s inequality to split the $\lvert k\pm\tilde{u}\rvert$. It follows that $\lVert U_{u}\hat{F}^{-1}\rVert_{r}\lesssim u^{\eta}$ when $r^{-1}>(2+\eta)/d$ and $\lVert\hat{F}_{\gamma}(k-\tilde{u})U_{u}\hat{F}^{-1}(k)\rVert_{q}\lesssim u^{\eta}$ when $q^{-1}>(\lvert\gamma\rvert+\eta)/d$.
Bound on $U_{u}(\hat{E}_{\gamma}/(\hat{A}\hat{F}))$. By hypothesis, $\lvert\gamma\rvert+\eta<\frac{1}{2}d+2+\rho$. We consider separately the two cases $\lvert\gamma\rvert<\sigma$ and $\lvert\gamma\rvert\geq\sigma$.
Case 1: $\lvert\gamma\rvert<\sigma$. By the quotient rule,
$\nabla_{j}(\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}})=\frac{\hat{E}_{\gamma+e_{j}}}{\hat{A}\hat{F}}-\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}}\frac{\hat{A}_{e_{j}}}{\hat{A}}-\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}}\frac{\hat{F}_{e_{j}}}{\hat{F}},$ (2.65)
which is in $L^{p}$ when $p^{-1}>(3-\sigma+\lvert\gamma\rvert)/d$ by (2.18) and Hölder’s inequality. Since $\lvert\gamma\rvert<\sigma$ and $d\geq 3$, we can take $p=1$, so $\hat{E}_{\gamma}/(\hat{A}\hat{F})$ is weakly differentiable. Lemma 2.9 then gives $\lVert U_{u}(\hat{E}_{\gamma}/(\hat{A}\hat{F}))\rVert_{q}\lesssim u^{\eta}$ for
$\frac{1}{q}=\frac{1}{p_{\eta}}=\frac{1}{p}-\frac{1-\eta}{d}>\frac{2-\sigma+\lvert\gamma\rvert+\eta}{d},$ (2.66)
which is the desired result.
Case 2: $\lvert\gamma\rvert\geq\sigma$. For $\sigma\leq\lvert\gamma\rvert<\frac{1}{2}d+2+\rho$, we apply (2.44) to obtain
$U_{u}(\frac{\hat{E}_{\gamma}}{\hat{A}\hat{F}})(k)$ $=\frac{\hat{E}_{\gamma}}{\hat{A}}(k+\tilde{u})U_{u}\hat{F}^{-1}(k)+\frac{U_{u}\hat{E}_{\gamma}(k)}{\hat{A}(k+\tilde{u})\hat{F}(k-\tilde{u})}+U_{u}\hat{A}^{-1}(k)\frac{\hat{E}_{\gamma}}{\hat{F}}(k-\tilde{u}).$ (2.67)
We show all three terms on the right-hand side have $L^{q}$ norm $u^{\eta}$ when $q^{-1}>(2-\sigma+\lvert\gamma\rvert+\eta)/d$. For the first term, we have already seen below (2.64) that $\lVert U_{u}\hat{F}^{-1}\rVert_{r}\lesssim u^{\eta}$ when $r^{-1}>(2+\eta)/d$. Since $\lvert\gamma\rvert\geq\sigma$, from (2.24) we have $\hat{E}_{\gamma}/\hat{A}\in L^{p}$ for $p^{-1}>(\lvert\gamma\rvert-\sigma)/d$. Hölder’s inequality then gives the desired $\lVert\frac{\hat{E}_{\gamma}}{\hat{A}}(k+\tilde{u})U_{u}\hat{F}^{-1}(k)\rVert_{q}\lesssim u^{\eta}$ for $q^{-1}=p^{-1}+r^{-1}>(2-\sigma+\lvert\gamma\rvert+\eta)/d$. The last term of (2.67) is analogous.
For the middle term of (2.67), we further divide into two subcases according to how $\lvert\gamma\rvert+\eta$ compares with $2+\sigma$. If $2+\sigma\leq\lvert\gamma\rvert+\eta<\frac{1}{2}d+2+\rho$, application of Lemma 2.10 to $x^{\gamma}E(x)$, which decays like $\lvert x\rvert^{-(d+2+\rho-\lvert\gamma\rvert)}$, gives
$\lVert U_{u}\hat{E}_{\gamma}\rVert_{r}\lesssim u^{\eta}\qquad(r^{-1}>\frac{\lvert\gamma\rvert+\eta-2-\sigma}{d}).$ (2.68)
The infrared bounds and Hölder’s inequality then yield $\lVert\frac{U_{u}\hat{E}_{\gamma}(k)}{\hat{A}(k+\tilde{u})\hat{F}(k-\tilde{u})}\rVert_{q}\lesssim u^{\eta}$ for $q^{-1}>(2-\sigma+\lvert\gamma\rvert+\eta)/d$. If instead $\lvert\gamma\rvert+\eta<2+\sigma$ then we apply Lemma 2.13 to obtain
$\lvert U_{u}\hat{E}_{\gamma}(k)\rvert\lesssim u^{\eta}(\lvert k+\tilde{u}\rvert^{2+\sigma-\eta-\lvert\gamma\rvert}+\lvert k-\tilde{u}\rvert^{2+\sigma-\eta-\lvert\gamma\rvert}).$ (2.69)
With the infrared bounds, this gives
$\Big{\lvert}\frac{U_{u}\hat{E}_{\gamma}(k)}{\hat{A}(k+\tilde{u})\hat{F}(k-\tilde{u})}\Big{\rvert}\lesssim u^{\eta}\left(\frac{1}{\lvert k+\tilde{u}\rvert^{\lvert\gamma\rvert-\sigma+\eta}\lvert k-\tilde{u}\rvert^{2}}+\frac{1}{\lvert k+\tilde{u}\rvert^{2}\lvert k-\tilde{u}\rvert^{\lvert\gamma\rvert-\sigma+\eta}}\right).$ (2.70)
Since $\lvert\gamma\rvert\geq\sigma$, splitting $\lvert k\pm\tilde{u}\rvert$ with Hölder’s inequality directly gives $\lVert\frac{U_{u}\hat{E}_{\gamma}(k)}{\hat{A}(k+\tilde{u})\hat{F}(k-\tilde{u})}\rVert_{q}\lesssim u^{\eta}$ for $q^{-1}>(2-\sigma+\lvert\gamma\rvert+\eta)/d$.
This completes the proof of Lemma 2.12. ∎
Appendix A Weak derivatives
We summarise here the properties of weak derivatives that we need. More can be found in [10, Chapter 5], [11, Chapter 4], or [28, Chapter 1]. We restrict our discussion to functions on the torus $\mathbb{T}^{d}$, which simplifies matters because $\mathbb{T}^{d}$ is compact and has no boundary. Let $C^{\infty}(\mathbb{T}^{d})$ denote the space of infinitely differentiable test functions $\phi:\mathbb{T}^{d}\to\mathbb{R}$. Since the torus is compact, these test functions have compact support. All integrals in this appendix are with respect to the normalised Lebesgue measure $dk/(2\pi)^{d}$ on $\mathbb{T}^{d}$.
Definition A.1.
Suppose $u,v\in L^{1}(\mathbb{T}^{d})$ and $\alpha$ is a multi-index. We say that $v$ is the $\alpha^{\rm th}$ weak partial derivative of $u$, written $\nabla^{\alpha}u=v$, if, for all test functions $\phi\in C^{\infty}(\mathbb{T}^{d})$,
$\int_{\mathbb{T}^{d}}u\nabla^{\alpha}\phi=(-1)^{\lvert\alpha\rvert}\int_{\mathbb{T}^{d}}v\phi.$ (A.1)
If $u$ is classically differentiable then (A.1) is the usual integration by parts formula, so weak derivatives are consistent with classical derivatives. The weak derivative is unique and satisfies standard calculus rules; see, e.g., [10, Sec. 5.2.3, Theorem 1] or [11, Theorem 4.4]. We make heavy use of the product and quotient rules, in a form that slightly generalises [11, Theorem 4.4], so we give the proofs here.
Lemma A.2 (Product rule).
Let $j\in\{1,\dots,d\}$. Suppose $f,g$ are weakly differentiable, and suppose that $f,\nabla_{j}f\in L^{p}$ and $g,\nabla_{j}g\in L^{q}$ with $\frac{1}{p}+\frac{1}{q}\leq 1$. Then the weak derivative $\nabla_{j}(fg)$ exists and is given by
$\nabla_{j}(fg)=(\nabla_{j}f)g+f(\nabla_{j}g).$ (A.2)
Proof.
By hypothesis and Hölder’s inequality, all of $fg$, $(\nabla_{j}f)g$, and $f(\nabla_{j}g)$ are in $L^{1}$. By convolving with a mollifier as in [11, Section 4.2.1], we take smooth approximations $f^{\varepsilon},g^{\varepsilon}\in C^{\infty}$ such that $f^{\varepsilon}\to f$ and $\nabla_{j}f^{\varepsilon}\to\nabla_{j}f$ in $L^{p}$, and $g^{\varepsilon}\to g$ and $\nabla_{j}g^{\varepsilon}\to\nabla_{j}g$ in $L^{q}$. Then, for $\phi\in C^{\infty}(\mathbb{T}^{d})$,
$\int_{\mathbb{T}^{d}}fg\nabla_{j}\phi=\lim_{\varepsilon\to 0}\int_{\mathbb{T}^{d}}f^{\varepsilon}g^{\varepsilon}\nabla_{j}\phi$ $=-\lim_{\varepsilon\to 0}\int_{\mathbb{T}^{d}}[(\nabla_{j}f^{\varepsilon})g^{\varepsilon}+f^{\varepsilon}(\nabla_{j}g^{\varepsilon})]\phi=-\int_{\mathbb{T}^{d}}[(\nabla_{j}f)g+f(\nabla_{j}g)]\phi,$ (A.3)
where the second equality uses classical integration by parts. This verifies the definition of the weak derivative in (A.1) and completes the proof. ∎
For the proof of the quotient rule, we use the ACL (absolutely continuous on lines) characterisation of weak derivative (see [28, Section 1.1.3] or [11, Theorem 4.21]), which is as follows. A function $u\in L^{1}(\mathbb{T}^{d})$ is weakly differentiable if and only if: (i) the restriction of $u$ to almost every line in the coordinate directions is absolutely continuous and (ii) the a.e.-defined classical partial derivatives are integrable over $\mathbb{T}^{d}$. In this case, the weak partial derivatives coincide with the classical partial derivatives.
Lemma A.3 (Quotient rule).
Let $j\in\{1,\dots,d\}$. Suppose that $f,g$ are weakly differentiable, that $g$ has countably many zeros, that $f/g\in L^{1}$, and that $[(\nabla_{j}f)g-f(\nabla_{j}g)]/g^{2}\in L^{1}$ for all $j$. Then $f/g$ is weakly differentiable, with
$\nabla_{j}\left(\frac{f}{g}\right)=\frac{(\nabla_{j}f)g-f(\nabla_{j}g)}{g^{2}}.$ (A.4)
Proof.
Since $f$ and $g$ are weakly differentiable, we know from the ACL characterisation that $f$ and $g$ are absolutely continuous on almost every line in the coordinate directions. Since $g$ has countably many zeros only, we can assume that $g$ is nonzero on these lines. Moreover, since $\mathbb{T}^{d}$ is compact, $g$ is in fact bounded away from 00 on these lines, so that $f/g$ is absolutely continuous on the lines. By the classical quotient rule, $f/g$ has classical partial derivatives almost everywhere, given by the right-hand side of (A.4), which are integrable by assumption. This verifies the ACL characterisation for $f/g$ and concludes the proof. ∎
The next lemma shows that weak derivatives preserve the classical fact that multiplication by a power corresponds to differentiation of the Fourier transform, a fact we rely on heavily. We write $\mathcal{F}[f]=\hat{f}$ for the $L^{2}$ Fourier transform of $f\in\ell^{2}(\mathbb{Z}^{d})$.
Lemma A.4.
Let $f:\mathbb{Z}^{d}\to\mathbb{R}$ and $\alpha$ be a multi-index. Suppose $x^{\alpha}f(x)\in\ell^{2}(\mathbb{Z}^{d})$. Then the $\alpha^{\rm th}$ weak partial derivative of $\hat{f}$ is given by
$\nabla^{\alpha}\hat{f}=\mathcal{F}[(ix)^{\alpha}f(x)].$ (A.5)
Proof.
Let $\phi\in C_{c}^{\infty}(\mathbb{T}^{d})$. By Parseval’s relation and integration by parts,
$\int_{\mathbb{T}^{d}}\mathcal{F}[(ix)^{\alpha}f(x)]\bar{\phi}$ $=\sum_{x\in\mathbb{Z}^{d}}(ix)^{\alpha}f(x)\overline{\mathcal{F}^{-1}[\phi]}$
$=(-1)^{\lvert\alpha\rvert}\sum_{x\in\mathbb{Z}^{d}}f(x)\overline{(ix)^{\alpha}\mathcal{F}^{-1}[\phi]}$
$=(-1)^{\lvert\alpha\rvert}\sum_{x\in\mathbb{Z}^{d}}f(x)\overline{\mathcal{F}^{-1}[\nabla^{\alpha}\phi]}=(-1)^{\lvert\alpha\rvert}\int_{\mathbb{T}^{d}}\hat{f}\,\overline{\nabla^{\alpha}\phi}.$ (A.6)
This verifies the definition of weak derivative (A.1), because $\phi$ is real. ∎
Acknowledgements
The work of both authors was supported in part by NSERC of Canada. We thank Takashi Hara and Emmanuel Michta for comments on a preliminary version. YL thanks Fanze Kong for discussions about weak derivatives.